OpenStudy (anonymous):

The number (1+sqrt 5)/2 is known as the golden ratio, and shows up in various natural solutions. The ancient Greeks used this number as a guide in their architectural designs for the ratio of width to height for building fronts. what is the conjugate expression for the golden ratio? my answer is 1-sqrt of 5 but the book says (1-sqrt 5)/2.. can someone explain it to me..

OpenStudy (experimentx):

(a+b)/b = b/a just put b/a = x and solve for it's two values.

OpenStudy (anonymous):

can explain further please?

OpenStudy (experimentx):

(a+b)/b = b/a <-- definition of golden ratio let b/a = golden ratio = x a/b+b/b = x 1/(b/a) + 1 = x 1/x + 1 = x <--- quadratic equation ... which has two roots. find two roots. that's all.

OpenStudy (phi):

the conjugate of 1+sqrt 5 is 1- sqrt 5 but they want the conjugate of (1+sqrt 5)/2

OpenStudy (anonymous):

that's it? lol i need to read carefully.. thank you very much:)